Three tacos cost a total of $1. 17. How many tacos can be purchased for $15.21?
Answer by nyc_function(2741)

Let t = tacos

3/t = 1.17/15.21

1.17t = 3(15.21)

1.17t = 45.63

t = 45.63/1.17

t = 39

You can buy 39 tacos.
NOTE: This is a typical proportion word problem. You will see tons of them in

math courses and math books. So, learn to identify them at first glance.